Suppose you wish to fabricate a uniform wire out o
Answer #1

R = 0.3 ohms

Mass of copper is 1.35 g = 0.00135 kg

Volume = Mass / Density

Density of copper = 8950 kg/m3

Volume = 0.00135/8950 = 1.50837*10-7 m3

Volume of a Cylinder = pi * radius2 * length

length = Volume / (pi * radius2) ............................................... 1

and Crosssectional Area of Cylinder = pi * radius2

R = p * L / A ............................................................... 2   

where p is resistivity (p of copper is 1.68*10-8 ohm-m)

Equating value of length in eqn 2 from 1

Resistance = (1.68*10-8 ohm-m)*(V/(pi*r2)) / (pi*r2)

(pi*r2)2 = 1.68*10-8 *1.50837*10-7/0.3 = 0.8446 *10-15

r4 = 8.567*10-16

r = 1.7108 *10-4 m, diameter = 2r = 3.4216*10-4 m = 0.3421 mm

L = 1.50837*10-7/pi* (1.7108 *10-4)2

L = 5.15 m